OpenStudy (anonymous):

the sum of three integers is two hundred eighty. The second number is ten less than twice the first number. The third number is four times the second number. Find the three numbers

OpenStudy (amistre64):

f +s +l = 280 s = 2f-10 l = 4(s) = 4(2f-10) = 8f -40 f + (2f -10) +(8f -40) = 280 11f -50 = 280 11f = 330 f = 30 ---------------- s = 2(30)-10 = 60 - 10 s = 50 ------------------------ l = 4(50) l = 200 ------------------- 30 + 50 +200 = 280 80 + 200 = 280 280 = 280

OpenStudy (anonymous):

are the three numbers 30, 60, and 80

OpenStudy (amistre64):

........ 30 50 200